# Forget What You Know about LLMs Evaluations - LLMs are Like a Chameleon

### v1

@@ Abstract
Large language models (LLMs) often appear to excel on public benchmarks, but these high scores may mask an overreliance on dataset-specific surface cues rather than true language understanding. We introduce the Chameleon Benchmark Overfit Detector (C-BOD), a meta-evaluation framework that systematically distorts benchmark prompts via a parametric transformation and detects overfitting of LLMs. By rephrasing inputs while preserving their semantic content and labels, C-BOD exposes whether a model’s performance is driven by memorized patterns. Evaluated on the MMLU benchmark using 26 leading LLMs, our method reveals an average performance degradation of 2.15% under modest perturbations, with 20 out of 26 models exhibiting statistically significant differences. Notably, models with higher baseline accuracy exhibit larger performance differences under perturbation, and larger LLMs tend to be more sensitive to rephrasings indicating that both cases may overrely on fixed prompt patterns. In contrast, the Llama family and models with lower baseline accuracy show insignificant degradation, suggesting reduced dependency on superficial cues. Moreover, C-BOD’s dataset- and model-agnostic design allows easy integration into training pipelines to promote more robust language understanding. Our findings challenge the community to look beyond leaderboard scores and prioritize resilience and generalization in LLM evaluation.
@@
Nurit Cohen-Inger1, Yehonatan Elisha2, Bracha Shapira1, Lior Rokach1, Seffi Cohen1 1Ben Gurion University  2Tel Aviv University
@@ 1 Introduction
Large Language Models (LLMs) have achieved impressive results on a wide range of NLP tasks Chang et al. (2024). Consequently, hundreds of benchmarks have been established to track progress and evaluate model capabilities Lu et al. (2024); Liang et al. (2022). However, the rapid proliferation of LLMs and the frequent use of public leaderboards raise concerns about the robustness of these evaluation practices Castillo-Bolado et al. (2024). Specifically, as benchmark data becomes more widely recognized, models may learn to exploit surface patterns or spurious correlations, rather than exhibit genuine language understanding. This issue can lead to deceptively high scores that do not reflect true progress.
In this paper, we examine whether LLMs rely excessively on benchmark-specific cues potentially overfitting to the patterns inherent in widely published evaluation benchmarks and explore systematic methods to detect and mitigate this behavior. In other words, are LLMs prone to overfitting on popular benchmarks, and what underlying factors contribute to this phenomenon?
To answer this question, we introduce Chameleon Benchmark Overfit Detector (C-BOD), a framework that reveals how heavily a model depends on the exact wording or structure of a test set. By introducing controlled textual distortions to benchmark prompts at varying intensities (defined by a distortion parameter $\mu$), as demonstrated in Figure 1, our method reveals whether high performance is built on superficial patterns. Notably, our framework requires only the evaluation set, without accessing the model’s training data or architecture. Unlike conventional leaderboards that solely track performance, our meta-evaluation framework acts as a safeguard ensuring that high scores do not stem from superficial memorization of benchmark cues.
@@
    Robust Overfitting Detection with Statistical Significance. Our framework computes the performance difference $\Delta_{\mu}$ between original and perturbed prompts and confirms its statistical significance, ensuring that observed differences indeed indicate overfitting rather than chance variations.
@@
    New Findings For LLM Community Our extensive analysis reveals new trends regarding how LLMs function with respect to their number of parameters and baseline performance.
@@
    Extensive Empirical Validation. We apply our method to multiple LLM families of various architectures and parameter sizes. Even modest textual distortions cause significant performance differences in most models, providing strong empirical evidence that overfitting is widespread.
@@
    Publicly Available Benchmarks and Code. We release rephrased versions of the widely used MMLU evaluation set under different distortion levels ($\mu$). These resources enable the community to adopt more robust, surface-invariant tests for reliable LLM assessment with our method reproducible code.
  5. 5.
    Blueprint for Iterative Overfit Mitigation. Beyond detection, these $\mu$-based rephrasings can be integrated into model training or fine-tuning pipelines. Regularly exposing models to diverse prompt variations helps reduce reliance on benchmark-specific phrasing, thus promoting more generalizable language understanding.
@@ 2 Related Work
2.1 Benchmark Evaluation and Overfitting
@@
LLMs have achieved impressive results on many benchmarks. This success has driven the development of comprehensive evaluation suites such as BIG-Bench Srivastava et al. (2022) and HELM Liang et al. (2022). MMLU benchmark set Hendrycks et al. (2020) evaluates question answering across 57 subjects—including STEM, humanities, and social sciences, while Zhang et al. (2024a) introduced 25 enterprise-focused datasets covering domains like finance, legal, cybersecurity, and climate sustainability for tasks such as classification, NER, and summarization. Another recent resource, JUDGE-BENCH Bavaresco et al. (2024), comprises 20 NLP datasets that assess models against human judgments. We focus on MMLU because of its widespread adoption and comprehensive domain coverage Wang et al. (2024), which makes it particularly effective for exposing overfitting to canonical prompt structures.
@@
While these benchmarks have been critical for comparing new models’ versions, recent studies warn that publicly released evaluation sets can become less reliable over time due to overexposure and memorization Yu et al. (2024); Chang et al. (2024). In some cases, LLMs learn superficial patterns specific to well-known datasets, boosting performance without reflecting genuine semantic or conceptual understanding. Kiela et al. (2021) further emphasize the need for continuously refreshing benchmarks to ensure real progress in language understanding. For example, OpenAI’s GPT models have shown steady improvement on MMLU: GPT-3 achieved approximately 43% accuracy in 2020 Brown et al. (2020), rising to nearly 70% with GPT-3.5 in 2022111https://cdn.openai.com/papers/GPT-4-Technical-Report.pdf, and reaching 86% with GPT-4 in 2023  [^2] https://cdn.openai.com/papers/GPT-4-Technical-Report.pdf.
@@
Memorization in LLMs has been widely studied Kiyomaru et al. (2024); Biderman et al. (2024), with larger models especially prone to retaining training data verbatim Carlini et al. (2022). This phenomenon can inflate performance metrics while obscuring genuine model capabilities. Moreover, several works highlight training-set contamination, where test samples appear exactly or as near-duplicates in the training data, as another crucial form of overfitting Deng et al. (2023); Yao et al. (2024), leading to overly optimistic performance estimates Yang et al. (2023).
2.2 Gap in Current Work
Researchers have introduced various methods to detect or mitigate training contamination: Finding the N-gram overlap (e.g., 13-grams or 50-character matches) between training and test data Brown et al. (2020); OpenAI (2023), though it can miss semantically equivalent rephrasing. Embedding similarity search Reimers (2019) that uses transformer-based embeddings to identify semantically close training-test pairs Lee et al. (2023). Decoding Matching probes the model by providing partial test prompts and measuring how likely it is to complete them exactly Li (2023) or completing missing words Deng et al. (2023). A recent study presented an overfit detection of editing knowledge to a LLM Zhang et al. (2025).
Although these studies have focused on detecting training data contamination or focusing on additional knowledge, they lack with addressing a critical issue: overfitting to benchmark-specific artifacts. In many cases, LLMs may never see the test data during training yet still learn to rely on superficial cues unique to a benchmark’s canonical format. Existing techniques such as n-gram overlap and embedding similarity fail to capture this subtle form of overfitting. In contrast, our approach explicitly quantifies the dependence of a model’s performance on the precise phrasing and structure of evaluation prompts, thereby filling this gap in current evaluation methodologies. By systematically applying a controllable distortion parameter to evaluation prompts, without requiring additional training or access to training data, our method shows how performance metrics degrade under textual perturbations, providing a robust means of diagnosing and mitigating broader overfitting behavior.
@@ 3 Method
Let $\mathcal{D}$ denote a benchmark dataset with N samples, and $\mathcal{E}$ a LLM to be evaluated with respect to a given performance function $\mathcal{M}$. Our goal is to detect whether $\mathcal{E}$ exhibits overfitting to $\mathcal{D}$. Figure 2 provides an overview of our proposed method, Chameleon Benchmark Overfit Detector (C-BOD). C-BOD employs a rephrasing transformation to generate a perturbed dataset from $\mathcal{D}$, evaluates on both the original and perturbed datasets, and applies a statistical test to assess whether performance discrepancies indicate overfitting. The following subsections detail each component of C-BOD.
Figure 2: High-level pipeline of our parametric approach. The original dataset $\mathcal{D}$ is passed through the distortion operator $T_{\mu}$ to form $\mathcal{D}_{\mu}$. Both sets are evaluated by a LLM, and differences in performance are used to quantify overfitting.
@@ 3.1 C-BOD rephrased dataset generation
To systematically introduce textual variations, C-BOD utilizes a rephrasing tool, denoted as ${T}$, which uses as a distortion operator to generate a perturbed dataset $\mathcal{D}_{\mu}$ from $\mathcal{D}$. This operator is parameterized by $\mu$ (temperature), which controls the extent of textual modification, ranging from low (e.g., 0.1 for minimal changes like synonym substitution) to moderate (e.g., 1.0 for rewording and sentence fragment reordering) and high (e.g., 1.5 for aggressive modifications such as question reformulation). We define:
@@ 3.2 Evaluating the Impact of Distortion
To assess the impact of distortion, we evaluate $\mathcal{E}$ using a performance function, $\mathcal{M}$. This function evaluates $\mathcal{E}$ based on a given ground truth or prompt $y_{i}$, considering two versions of an input: the original $x_{i}\in\mathcal{D}$ and the perturbed version $x_{i}^{\prime}\in\mathcal{D}_{\mu}$, where $i$ denotes the index of a sample in the dataset. Specifically, $\mathcal{M}$ is a boolean function that takes as input the model $\mathcal{E}$ and two data pairs, $(x_{i},y_{i})$ and $(x_{i}^{\prime},y_{i})$, and returns whether the model performs better on the original input than on the perturbed one. The function is formulated as follows:
$\mathcal{M}(\mathcal{E},(x_{i},y_{i}),(x^{\prime}_{i},y_{i})=\begin{cases}1,&\text{if }P(\mathcal{E},x_{i},y_{i})\\ &\quad>P(\mathcal{E},(x^{\prime}_{i},y_{i}),\\ 0,&\text{otherwise.}\end{cases}$
@@
$\Delta_{\mu},b=\sum_{i=0}^{N}\mathcal{M}(\mathcal{E},(x_{i},y_{i}),(x_{i}^{\prime},y_{i}))$ (1)
@@ 3.3 Statistical Validation
2:$\mathcal{D}$: Original benchmark dataset of size $N$,
$\mathcal{E}$: LLM,
4:$\mu$: Distortion parameter,
5:$T_{\mu}$: Transformation operator,
$\mathcal{M}$: Performance function (returns 1 if the first input is better, 0 otherwise),
7:$\alpha$: Significance level.
@@
17:$\chi^{2}\leftarrow\dfrac{(b-c)^{2}}{b+c}$
18:$p\leftarrow\text{p-value}(\chi^{2},\text{df}=1)$
@@
28:return $Overfit\_Flag$
@@ 4.1 Dataset and Rephrasing Process
Our experiments use the MMLU benchmark Hendrycks et al. (2020), which comprises multiple-choice questions spanning 57 subjects. The broad coverage and public availability of MMLU make it an ideal candidate for assessing general knowledge and the degree to which LLMs overfit canonical prompt formats. The MMLU dataset is distributed under the MIT License, which allows for free use, modification, and distribution as long as the original copyright notice and license terms are maintained. We generate a perturbed versions of the original dataset to probe overfitting, following the methodology described in Section 3. We used DeepSeek 3 to create the transformed version of each question. We generate the perturbed dataset using $\mu=1.0$ (the default temperature parameter).
These perturbations include synonym substitutions, sentence reordering, and the insertion of distractor phrases, while preserving the original semantic meaning and correct answers. Automated formatting checks and manual audits (performed on approximately 10% of the samples) ensure that the integrity of the questions is maintained. For example, an original question:“The coronal suture joins the?” is rephrased as:“Which bones does the coronal suture connect?”. The perturbed dataset, denoted by $\mathcal{D}_{\mu}$, is released alongside our code for reproducibility  [^3] https://github.com/SeffiCohen/CBOD.

### v2

@@ Abstract
Large language models (LLMs) often appear to excel on public benchmarks, but these high scores may mask an overreliance on dataset-specific surface cues rather than true language understanding. We introduce the Chameleon Benchmark Overfit Detector (C-BOD), a meta-evaluation framework designed to reveal such overfitting. C-BOD systematically rephrases benchmark inputs via a parameterized transformation that preserves semantic content and labels, enabling the detection of performance degradation indicative of superficial pattern reliance. We conduct extensive experiments across two datasets, three rephrasing models, and multiple distortion levels, evaluating 32 state-of-the-art LLMs. On the MMLU benchmark, C-BOD reveals an average performance drop of 2.75% under modest rephrasings, with over 80% of models exhibiting statistically significant differences. Notably, higher-performing models and larger LLMs tend to show greater sensitivity, suggesting a deeper dependence on benchmark-specific phrasing. Due to its dataset and model-agnostic design, C-BOD can be easily integrated into evaluation pipelines and offers a promising foundation for overfitting mitigation strategies. Our findings challenge the community to look beyond leaderboard scores and prioritize resilience and generalization in LLM evaluation. Our code and benchmark datasets are available at: https://github.com/nuritci/cbod
@@
Nurit Cohen-Inger1, Yehonatan Elisha2, Bracha Shapira1, Lior Rokach1, Seffi Cohen3 1Ben Gurion University  2Tel Aviv University  3Harvard University
@@ 1 Introduction
Large Language Models (LLMs) have achieved impressive results on a wide range of NLP tasks Chang et al. (2024). Consequently, hundreds of benchmarks have been established to track progress and evaluate model capabilities Lu et al. (2024); Liang et al. (2022). However, the rapid proliferation of LLMs and the frequent use of public leaderboards raise concerns about the robustness of these evaluation practices Castillo-Bolado et al. (2024). Specifically, as benchmark data becomes more widely recognized, models may learn to exploit surface patterns or spurious correlations, rather than exhibit genuine language understanding. This issue can lead to deceptively high scores that do not reflect true progress. In this paper, we examine whether LLMs rely excessively on benchmark-specific cues potentially overfitting to the patterns inherent in widely published evaluation benchmarks and explore systematic methods to detect and mitigate this behavior. In other words, are LLMs prone to overfitting on popular benchmarks, and what underlying factors contribute to this phenomenon? To answer this question, we introduce the Chameleon Benchmark Overfit Detector (C-BOD). This framework reveals how heavily a model depends on the exact wording or structure of a test set. By introducing controlled textual distortions to benchmark prompts at varying intensities (defined by a distortion parameter $\mu$), as demonstrated in Figure 1, our method exposes whether strong performance derives from reliance on superficial patterns. Notably, our framework requires only the evaluation set, without accessing the model’s training data or architecture. Unlike conventional leaderboards that solely track performance, our meta-evaluation framework acts as a safeguard ensuring that high scores do not stem from superficial memorization of benchmark cues.
@@
    A Robust Framework for Detecting Benchmark Overfitting. We present a framework that computes the performance difference $\Delta_{\mu}$ between original and perturbed prompts and confirms its statistical significance, ensuring that observed differences indeed indicate overfitting rather than chance variations.
@@
    New Insights into LLM Behavior. Our analysis reveals that larger models and those with higher baseline performance are often more sensitive to perturbations, suggesting a deeper reliance on benchmark-specific phrasing.
@@
    Extensive Empirical Validation. We apply our method to a diverse collection of 32 leading LLMs from various families, architectures, and parameter sizes. Utilizing modest textual distortions generated by three different rephrasing models for enhanced robustness, our analysis reveals statistically significant performance degradation in over 80% of the evaluated models, providing strong empirical evidence of widespread benchmark overfitting. Similar trends were observed in our evaluation on the GPQA dataset, in which we also show an ablation over different distortion levels ($\mu$).
@@
    Open Resources for the Community. To facilitate further research and promote robust evaluation, we publicly release the rephrased versions of the widely used MMLU and GPQA evaluation sets, along with our reproducible code. These resources enable the community to readily adopt more robust, surface-invariant tests for reliable LLM assessment and provide a foundation for developing mechanisms to mitigate benchmark overfitting.
@@ 2 Related Work
2.1 Benchmark Datasets and Evaluation Suites
@@
LLMs have achieved impressive results on many benchmarks. This success has driven the development of comprehensive evaluation suites such as BIG-Bench Srivastava et al. (2022) and HELM Liang et al. (2022). The MMLU benchmark set Hendrycks et al. (2020) evaluates question answering across 57 subjects, including STEM, humanities, and social sciences, while Zhang et al. (2024) introduced 25 enterprise-focused datasets covering domains like finance, legal, cybersecurity, and climate sustainability for tasks such as classification, NER, and summarization. Another recent resource, JUDGE-BENCH Bavaresco et al. (2024), comprises 20 NLP datasets that assess models against human judgments. GPQA benchmark Rein et al. (2024) is tasked with evaluating reasoning models. We focus on MMLU and GPQA because of their widespread adoption [^1] https://klu.ai/glossary/gpqa-eval and comprehensive domain coverage Wang et al. (2024).
2.2 Overfitting in LLMs
@@
While these benchmarks have been critical for comparing new models’ versions, recent studies warn that publicly released evaluation sets can become less reliable over time due to overexposure and memorization Yu et al. (2024); Chang et al. (2024). In some cases, LLMs learn superficial patterns specific to well-known datasets, boosting performance without reflecting genuine semantic or conceptual understanding. Kiela et al. (2021) further emphasizes the need for continuously refreshing benchmarks to ensure real progress in language understanding. For example, OpenAI’s GPT models have shown steady improvement on MMLU: GPT-3 achieved approximately 43% accuracy in 2020 Brown et al. (2020), rising to nearly 70% with GPT-3.5 in 2022, and reaching 86% with GPT-4 in 2023 Koubaa (2023).
@@
Memorization in LLMs has been widely studied Kiyomaru et al. (2024); Biderman et al. (2024), with larger models especially prone to retaining training data verbatim Carlini et al. (2022). This phenomenon can inflate performance metrics while obscuring genuine model capabilities. Moreover, several works highlight training-set contamination, where test samples appear exactly or as near-duplicates in the training data, as another crucial form of overfitting Deng et al. (2023); Yao et al. (2024), leading to overly optimistic performance estimates Yang et al. (2023). Training Data Contamination refers to the presence of test data, or near duplicates, within the training set Deng et al. (2023); Yao et al. (2024). Contamination renders evaluation unreliable, as the model has effectively already seen the "test" data, leading to overly optimistic performance estimates Yang et al. (2023). Benchmark/Prompt Structure Overfitting is a more subtle form that arises when LLMs learn to exploit superficial cues or patterns specific to a benchmark’s format or the structure of evaluation prompts, even without memorizing the exact content Yu et al. (2024). This can lead to overestimated generalization ability, as the model’s performance becomes dependent on the specific benchmark artifacts rather than true language understanding. This type of overfitting is the focus of our work.
2.3 Addressing the Gap
Current methods largely overlook the crucial problem of overfitting to benchmark-specific artifacts, which can significantly misrepresent an LLM’s true capabilities and hinder the development of robust and generalizable models. Our work addresses this gap by introducing a novel method to quantify an LLM’s reliance on benchmark prompt structure. We systematically apply controlled distortions to evaluation prompts, for example, by replacing synonyms or altering word order and measure the resulting performance degradation. This approach, which does not require access to training data, provides a direct measure of vulnerability to prompt structure and a robust means of diagnosing and mitigating this critical form of overfitting.
@@ 3 Method
Let $\mathcal{D}$ denote a benchmark dataset with N samples, and $\mathcal{E}$ an LLM to be evaluated with respect to a given performance function $\mathcal{M}$. Our goal is to detect whether $\mathcal{E}$ exhibits overfitting to $\mathcal{D}$. Figure 2 provides an overview of our proposed method, Chameleon Benchmark Overfit Detector (C-BOD). C-BOD employs a rephrasing transformation to generate a perturbed dataset from $\mathcal{D}$, evaluates on both the original and perturbed datasets, and applies a statistical test to assess whether performance discrepancies indicate overfitting. The following subsections detail each component of C-BOD.
Figure 2: High-level pipeline of our parametric approach. The original dataset $\mathcal{D}$ is passed through the distortion operator $T_{\mu}$ to form $\mathcal{D}_{\mu}$. Both sets are evaluated by an LLM, and differences in performance are used to statistically quantify overfitting.
@@ 3.1 C-BOD rephrased dataset generation
To systematically introduce textual variations, C-BOD utilizes a rephrasing tool, denoted as T{T}, which uses as a distortion operator to generate a perturbed dataset $\mathcal{D}_{\mu}$ from $\mathcal{D}$. This operator is parameterized by $\mu$, which controls the extent of textual modification, ranging from low (e.g., 0.1 for minimal changes like synonym substitution) to moderate (e.g., 0.5 for rewording and sentence fragment reordering) and high (e.g., 1.5 for aggressive modifications such as question reformulation). Specifically, for an LLM, $\mu$ corresponds directly to the temperature used during generation, influencing the degree of variation in rephrased outputs. We define:
@@ 3.2 Evaluating the Impact of Distortion
To assess the impact of distortion, we evaluate $\mathcal{E}$ using a performance function, $\mathcal{M}$. This function evaluates $\mathcal{E}$ based on a given ground truth $y_{i}$, considering two versions of an input: the original $x_{i}\in\mathcal{D}$ and the perturbed version $x_{i}^{\prime}\in\mathcal{D}_{\mu}$, where $i$ denotes the index of a sample in the dataset. Specifically, $\mathcal{M}$ is a boolean function that takes as input the model $\mathcal{E}$ and two data pairs, $(x_{i},y_{i})$ and $(x_{i}^{\prime},y_{i})$, and returns whether the model performs better on the original input than on the perturbed one. The function is formulated as follows:
$\mathcal{M}(\mathcal{E},(x_{i},y_{i}),(x^{\prime}_{i},y_{i}))=\begin{cases}1,&\text{if }P(\mathcal{E},x_{i},y_{i})\\ &\quad>P(\mathcal{E},x^{\prime}_{i},y_{i}),\\ 0,&\text{otherwise.}\end{cases}$
@@
$\Delta_{\mu}=b=\sum_{i=0}^{N}\mathcal{M}(\mathcal{E},(x_{i},y_{i}),(x_{i}^{\prime},y_{i}))$ (1)
@@ 3.3 Statistical Validation
[…]
@@
[…]
@@
28:return Overfit_FlagOverfit\_Flag
@@ 4.1 Dataset and Rephrasing Process
Our experiments used two leading benchmark datasets: (1) MMLU Hendrycks et al. (2020): This benchmark spans 57 subjects, comprising 14,079 test samples and 1,540 validation samples. Its broad coverage makes it a standard choice for evaluating general knowledge and assessing overfitting to canonical prompt formats. It is distributed under the MIT License, allowing free use and modification. (2) GPQA Rein et al. (2024): Introduced in 2024, this smaller yet challenging benchmark includes 546 multi-step reasoning samples designed to test logical inference. Its complex, less common question structures help reduce the risk of training data contamination, making it a valuable complement to MMLU for overfitting analysis. Detailed results and analysis on GPQA are provided in Appendix C.
We generate a perturbed version of the original dataset to probe overfitting, following the methodology described in Section 3. We used DeepSeek 3 to create the transformed version of each question and generate the perturbed dataset $\mathcal{D}_{1.0}$ using $\mu=1.0$ (the default temperature parameter for DeepSeek), and Claude 3.5 Haiku to generate the perturbed dataset $\mathcal{D}_{0.5}$ using $\mu=0.5$ (the default temperature parameter for Claude). For the GPQA we used GPT-4o-mini to generate the perturbed datasets $\mathcal{D}_{0.5}$, $\mathcal{D}_{1.0}$, $\mathcal{D}_{1.5}$, using $\mu\in\{0.5,1.0,1.5\}$.
These perturbations include synonym substitutions, sentence reordering, and the insertion of distractor phrases, while preserving the original semantic meaning and correct answers. The perturbed datasets, denoted by $\mathcal{D}_{\mu}$, is released alongside our code for reproducibility.
4.1.1 Evaluation of Rephrasing Quality
To ensure the quality of rephrasing in the C-BOD framework, we implemented a multi-step evaluation approach to maintain the semantic integrity of the original prompts and avoid confounding overfitting assessments. This process included three main validation steps: (1) Cosine Similarity Analysis, (2) Semantic Equivalence Verification, and (3) Iterative Human Audits. In what follows, we briefly describe these steps.
Cosine Similarity Analysis
We measured semantic alignment using Sentence-BERT Reimers and Gurevych (2019), confirming high alignment across perturbations:
  • MMLU 0.5 (Claude):
    Mean = 0.829, Median = 0.864.
  • MMLU 1.0 (DeepSeek):
    Mean = 0.883, Median = 0.922.
  • GPQA 0.5 (GPT):
    Mean = 0.954, Median = 0.969.
  • GPQA 1.0:
    Mean = 0.947, Median = 0.967.
  • GPQA 1.5:
    Mean = 0.942, Median = 0.961.
Semantic Equivalence Verification
In addition to cosine similarity, we employed a reasoning model GPT-o3 to independently verify the retention of original intent. Approximately 1-2% of the rephrasings exhibited semantic errors or insufficient similarity, requiring further manual correction and refinement before inclusion in the final evaluation set. This additional verification reduced the risk of subtle semantic drift, ensuring a high-quality perturbation set.
Iterative Human Audits
To ensure the accuracy and semantic consistency of the perturbed datasets, we conducted a comprehensive manual audit as the final validation step. This process specifically targeted prompts with a cosine similarity score below 0.7 or those that received a negative judgment from the automated evaluation. These prompts were iteratively refined through API adjustments until they met the desired fidelity thresholds. Overall, this approach manually covered approximately 20% of the MMLU dataset and 50% of the GPQA dataset, supplementing the earlier automated checks that covered 100% of both datasets. This multi-step approach significantly enhances the reliability of our overfitting assessments by ensuring precise, controlled textual variations.